Solve for X
log3^4 - log3^x =2

Answer by lwsshak3(11628)

Solve for X
log3^4 - log3^x =2
**
log(3^4/3^x)=2
log(3^(4-x)=2
(4-x)log3=2
4-x=2/log3=4.1918
x=4-4.1918=-.1918
..
Check:
log3^4 - log3^x
log81-log(3^-.1918)
1.9085-log.81
1.9085-(-.0915)=2